OpenStudy (anonymous):

Prove cosh(2x) = cosh^(2)x + sin^(2)x

OpenStudy (anonymous):

its just going back to the definitions

OpenStudy (anonymous):

\[\cosh (nx) = \frac{1}{2} ( e^{nx} + e^{-nx} ) \]

OpenStudy (anonymous):

\[\sinh(nx) = \frac{1}{2}(e^{nx} - e^{-nx} )\]

OpenStudy (anonymous):

\[(\sinh(x))^2 + (\cosh(x))^2 = \frac{1}{4} ( e^{x} + e^{-x} ) ^2 + \frac{1}{4}(e^{x} - e^{-x})^2\]

OpenStudy (anonymous):

then expand , you can do the rest , its pretty easy

OpenStudy (anonymous):

all you need to know is (e^(x))^2 = e^(2x)

OpenStudy (anonymous):

then alot of things will cancel and you will get \[\frac{1}{2}(e^{2x} +e^{-2x} ) \]

OpenStudy (anonymous):

which is cosh(2x) by definition